OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2 (x + 4)?

OpenStudy (ny,ny):

slope intercept form --> y=mx+b

OpenStudy (ny,ny):

get the y by itself on one side and distribute wgats in the parentheses to get m and b.

OpenStudy (ny,ny):

whats*

OpenStudy (anonymous):

so it would be y=–5/2 (x + 11)?

OpenStudy (anonymous):

how do i get rid of the x thoe?

OpenStudy (ny,ny):

hmmm, i got y=-5/2x-3

OpenStudy (ny,ny):

dont get rid of the x. it is part of slop intercept form.

OpenStudy (anonymous):

lol okay

OpenStudy (ny,ny):

y-7=-5/2(x+4) y-7=-5/2x-10 <-- distribute whats in parentheses +7 +7 ------------- y=-5/2x-3

OpenStudy (ny,ny):

see what i did?

OpenStudy (anonymous):

yes yes i did

OpenStudy (ny,ny):

good.

OpenStudy (anonymous):

thanks !

OpenStudy (ny,ny):

welcome :D